OpenStudy (anonymous):

What is the sum of series if nth term is (2n+1)^2

OpenStudy (anonymous):

It is \[\Sigma n\] where n is nth term ie,(\[\Sigma (2n+1)^2=\Sigma (4n^2+4n+1)=4\Sigma (n^2)+4\Sigma (n)+\Sigma (1)\] We have \[\Sigma (n^2)= n(n+1)(2n+1)/6\] \[\Sigma (n)=n(n+1)/2\] \[ \Sigma(1)=n\] Substituting these in above expression and simplify to get the answer.

OpenStudy (anonymous):

Answer is \[(4n^3+12n^2+11n)/6\]